OpenStudy (anonymous):

0.25r - 0.125 + 0.5r = 0.5 + r

OpenStudy (anonymous):

0.25r - 0.125 + 0.5r = 0.5 + r 0.25r + 0.5r - 0.125 = 0.5 + r 0.75r - 0.125 = 0.5 + r -0.125 - 0.5 = 0.75r + r -0.625 = 1.75r -0.625/1.75 = r -0.35714 = r